# Fusion of Movement and Naïve Predictions for Point Forecasting in Univariate Random Walks

### v2

@@ 4.1 Experiment 1: Simulated data
Figure 1 shows a simulated random walk series of 2,500 steps generated for evaluating the proposed fusion method. Each step of the random walk was drawn from a normal distribution with a mean of zero and a standard deviation of one, reflecting typical stochastic fluctuations. The cumulative sum of these steps constitutes a random walk path. To maintain all the values within a positive range, an initial positive offset of 50 units was applied to the entire series. The last 200 data points were selected as the test set, while the remaining data points were used to calculate $\bar{\epsilon}$, a constant value representing the mean of the absolute residuals of the naïve forecast derived from the historical data.
@@
Figure 2 illustrates the fusion forecasts using optimal weights across various movement prediction accuracies. When $ACC^{\text{movement}}$ slightly exceeds 0.50, such as in the first plot where $ACC^{\text{movement}}=0.55$, the fusion forecast shows only minor deviations from the naïve forecast and still exhibits a noticeable “shift” from the previous actual values. As $ACC^{\text{movement}}$ increases, this “shift” progressively diminishes and vanishes when $ACC^{\text{movement}}=1$.
To evaluate the significance of the differences observed between the fusion forecasts and the naïve forecast, the fusion process was repeated 30 times at each accuracy level, with all trials using random movement predictions of the same accuracy. Therefore, 30 independent MAE and MSE measures were generated at each level of $ACC^{\text{movement}}$, as shown in Figure 3. These box plots compare the MAEs and MSEs of the fusion forecasts to those of the naïve forecast across different levels of $ACC^{\text{movement}}$, indicating that the fusion forecasts generally achieved lower MAEs and MSEs than did the naïve forecast. Notably, most trials at low levels of $ACC^{\text{movement}}$ still outperformed the naïve forecast, such as when $ACC^{\text{movement}}=0.55$ and $ACC^{\text{movement}}=0.60$. Subsequently, the one-sample Wilcoxon signed-rank test was conducted to determine whether, at each accuracy level, the median value of the differences between the MAEs or MSEs of the fusion forecasts and those of the naïve forecast significantly deviated from zero. The MAEs and MSEs of the naïve and fusion forecasts, along with the improvements achieved by the fusion method over the naïve forecast and the results of Wilcoxon tests are summarized in Table 1. Notably, the largest p value, 0.0024, derived when $ACC^{\text{movement}}=0.55$, still indicates a statistically significant enhancement of the fusion forecast over the naïve forecast, confirming that the observed improvements are not due to chance and underscoring the effectiveness of the fusion approach even at relatively low levels of $ACC^{\text{movement}}$. As $ACC^{\text{movement}}$ increases, the disparity between the naïve and fusion forecasts increases significantly.
@@
Prediction MAE MSE
Range Improvement p value Range Improvement p value
Naïve 0.821 - - 1.044 - -
Acc = 0.55, c = 0.1 0.815 $\pm$ 0.006 0.005 0.0001 1.037 $\pm$ 0.012 0.008 0.0024
Acc = 0.60, c = 0.2 0.804 $\pm$ 0.010 0.016 0.0000 1.022 $\pm$ 0.021 0.022 0.0000
Acc = 0.65, c = 0.3 0.777 $\pm$ 0.017 0.043 0.0000 0.983 $\pm$ 0.037 0.062 0.0000
Acc = 0.70, c = 0.4 0.742 $\pm$ 0.021 0.078 0.0000 0.939 $\pm$ 0.042 0.105 0.0000
Acc = 0.75, c = 0.5 0.710 $\pm$ 0.024 0.111 0.0000 0.879 $\pm$ 0.043 0.165 0.0000
Acc = 0.80, c = 0.6 0.667 $\pm$ 0.027 0.154 0.0000 0.803 $\pm$ 0.050 0.241 0.0000
Acc = 0.85, c = 0.7 0.628 $\pm$ 0.013 0.192 0.0000 0.732 $\pm$ 0.041 0.313 0.0000
Acc = 0.90, c = 0.8 0.578 $\pm$ 0.027 0.243 0.0000 0.617 $\pm$ 0.069 0.427 0.0000
Acc = 0.95, c = 0.9 0.528 $\pm$ 0.017 0.293 0.0000 0.492 $\pm$ 0.041 0.553 0.0000
Acc = 1, c = 1 0.482 $\pm$ 0.000 0.339 0.0000 0.371 $\pm$ 0.000 0.673 0.0000
Figure 4 illustrates the performance of the fusion forecasts across various accuracy levels using different weights. As shown in the first four plots, under a smaller value of $c$, such as 0.01 or 0.1, the fusion forecast generally outperforms the naïve forecast, albeit at the expense of potential gains from higher $c$. Conversely, setting a larger $c$ could lead to inferior performance to the naïve forecast if the actual $ACC^{\text{movement}}$ is poor.
@@ 4.2 Experiment 2: Real-world Data
Table 3 presents the augmented Dickey–Fuller (ADF) test results for the two real-world time series. The ADF test can identify the presence of a unit root, assessing whether the S&P 500 and BTC price series exhibit random walk characteristics. For the S&P 500, the ADF statistic was 0.416927, with a p value of 0.982117. Given the critical values at the 1% (-3.433), 5% (-2.863), and 10% (-2.567) significance levels, the ADF statistic does not surpass these thresholds, and the high p value further supports the nonrejection of the null hypothesis. This suggests that the S&P 500 series likely follows a random walk, in which future values primarily reflect random fluctuations from previous values. Similarly, the BTC price series showed an ADF statistic of -1.449327 and a p value of 0.558413, with critical values not exceeded. This result indicates that the BTC price series is nonstationary and may also follow a random walk, implying that historical BTC prices are not indicative of future prices and that the series shows no tendency to revert to a long-term trend or mean.
@@
Although the accuracy of financial time series movement forecasts can range between 0.55 and 0.8, this study opted to not employ complex models that integrate external data sources, which are resource intensive. Instead, this study utilized the nu-support vector machine (nu-SVM), a reliable binary classification model with the proportion of support vectors tuned through a parameter nu Bishop and Nasrabadi (2006). This choice allows us to focus on testing the robustness of the fusion method under real movement prediction with moderate accuracy rather than pursuing higher accuracy through more sophisticated models. In addition, multiple trials with varying prediction accuracies were deemed unnecessary, as previous simulations have adequately demonstrated the performance of the fusion method across different accuracy levels. In addition, the details regarding the feature extraction process, which leverages a range of technical indicators and market statistics as model inputs, are outlined in Table LABEL:table:A1 in the Appendix. The nu-SVM prediction accuracy on the validation set was used as an estimation of $ACC^{\text{movement}}$ for setting the value of $c$ based on (11).
@@
Figure 6 and Figure 7 present the forecast comparisons of different methods for the S&P 500 and BTC series, respectively. Because the movement prediction accuracy on the validation set is approximately 0.55 for both time series, the weight $c$ based on (11) is approximately 0.1. Consistent with the theoretical analysis presented in Section 3.2, this small weight $c$ leads to a small discrepancy between the naïve and fusion forecasts. Complex models such as MLP and LSTM differ significantly from fusion, naïve, and ARIMA forecasts because they presuppose a mapping from historical sequences to future values, which conflicts with the fundamental definition of a random walk, where the future value depends solely on the sum of the current value and a random increment. The assumed patterns and dependencies in the data obtained by these models may not exist in true random walks.
Figure 6: Comparison of the S&P 500 forecast results of different methods.
Figure 7: Comparison of the BTC forecast results of different methods.
Table 5 summarizes the model performances in terms of the MAE and MSE values. The fusion method demonstrated the best overall performance among the models tested, achieving lower MAE and MSE values than did the naïve, ARIMA, and linear regression methods for both prediction tasks. Specifically, the fusion method yielded an MAE of 16.19 and an MSE of 490.91 for the S&P 500 prediction and an MAE of 431.94 with an MSE of 449611.83 for the BTC prediction, suggesting that the proposed method has a robust ability to make forecasts with reduced errors compared to other models. In contrast, both the MLP and LSTM models underperformed compared with the more traditional approaches, showing particularly high forecasting errors for both prediction tasks.
@@
MAE MSE MAE MSE
Fusion 16.187902 490.912977 431.940881 449611.828348
Naïve 16.277100 499.378612 433.766817 456875.143422
ARIMA 16.338302 499.739638 431.050006 456466.453451
Linear Regression 16.236584 497.368764 463.574310 486329.417207
MLP 21.615049 775.427551 496.776367 486248.250000
LSTM 21.148510 690.336853 457.213593 503120.968750
@@ 5 Discussion
This section discusses the implications of the findings based on theoretical analysis and experimental results and the potential limitations of this study. The empirical results show that when the $ACC^{\text{movement}}$ is approximately 0.55, the proposed fusion method outperforms the naïve forecast, which is consistent with the theoretical expectation that an accuracy above 0.5 is necessary. Interestingly, the accuracy of financial time series movement prediction can reach as high as 0.8 in some scenarios Bustos and Pomares-Quimbaya (2020), indicating that high movement prediction for real-world random walks is possible and can potentially lead to fusion results better than the examples shown in this study. Nonetheless, in real forecasting scenarios, $ACC^{\text{movement}}$ can only be estimated through model performance on the training or validation set, and such estimations can overestimate or underestimate the true $ACC^{\text{movement}}$. Whether $ACC^{\text{movement}}$ on the test set can consistently meet the accuracy threshold for improvement remains a challenge.
From the decision fusion perspective, integrating diverse predictions is advantageous because each model contributes unique insights and enhances predictive capabilities Bishop and Nasrabadi (2006). Previously, effectively combining predictions from various models that produce different types of data, such as binary and continuous outputs, has presented a significant challenge. The fusion method proposed in this study is an effective approach for converting binary data into continuous data, thereby enabling the fusion of diverse data types. This conversion facilitates an efficient integration process, allowing for the comprehensive use of available forecasts to improve decision making in real forecasting scenarios.
Notably, the theoretical improvement described in (10) differs from the empirical improvement observed in the experiments. Various factors could explain these discrepancies. The dataset used for experiments may not be large enough to meet the assumptions made in the theoretical analysis. In addition, while the definition of a random walk adopted in this study assumes i.i.d. increments for each time step, real-world data can be drifted random walks and are subject to measurement errors. Moreover, the random increments are supposed to be the same for all future time steps, which may not always provide a reliable estimation owing to the fluctuating nature of real-world data.
@@ 6 Conclusion
This study has made substantial contributions to the field of time series forecasting and information fusion by introducing a novel fusion method specifically designed for predicting random walk series future values. Key findings reveal that the fusion method can outperform naïve forecasts even with a movement prediction accuracy of approximately 0.55. This improvement is achieved without consuming large amounts of computational resources and significantly exceeds that of many traditional forecasting methods, which often fail to surpass naïve forecasts. The implications of these results are profound, suggesting that the fusion method offers a potent approach for enhancing the forecasting of random walk series by effectively utilizing movement predictions.
@@
There are several avenues for further research that build upon the foundations laid by this study. First, although this research focuses on random walk forecasting, future work could expand the forecasting philosophy presented in this paper to encompass other types of time series data, potentially broadening its applicability. Second, given the dependency of the proposed method on movement predictions, which are still challenging in various areas, there is a pressing need to develop more advanced movement prediction techniques for random walk series. Such advancements, along with the fusion method proposed in this study, could significantly enhance point forecasting in random walks.
Moreover, the current model utilizes a fixed increment for simplicity; however, adapting this increment based on the availability of new data or real-time estimation of the movement prediction accuracy may provide more dynamic and accurate forecasts. Investigating whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research. This adaptive approach could lead to more nuanced and responsive forecasting models, potentially transforming predictive analytics in various areas.

### v3

@@ 4.1 Experiment 1: Simulated data
Figure 1 shows a simulated random walk series of 2,500 steps generated for evaluating the proposed fusion method. Each step of the random walk was drawn from a normal distribution with a mean of zero and a standard deviation of one, reflecting typical stochastic fluctuations. The cumulative sum of these steps constitutes a random walk path. To maintain all the values within a positive range, an initial positive offset of 50 units was applied to the entire series. The last 200 data points were selected as the test set, while the remaining data points were used to calculate the value of $\bar{\epsilon}$.
@@
To evaluate the significance of the differences observed between the fusion forecasts and the naïve forecasts, the fusion process was repeated 100 times at each accuracy level, with all trials using random movement predictions of the same accuracy. Therefore, 100 independent MSE and MAE measures were generated at each $ACC^{\text{movement}}$ level, as shown in Figure 2. Each box plot represents a set of 100 independent MSE or MAE measures at a certain $ACC^{\text{movement}}$ level. Although variation was observed among each set, the fusion result generally improved as $ACC^{\text{movement}}$ increased, and except when $ACC^{\text{movement}}=0.5$, the fusion forecasts generally achieved lower MSEs and MAEs than did the naïve forecasts. When the movement accuracy is equal to 0.5, the movement prediction is essentially equivalent to a random guess. Under these conditions, the proposed fusion method does not offer any improvement over the naïve prediction. Consequently, the optimal prediction in this scenario is simply the naïve prediction itself. Subsequently, the one-sample Wilcoxon signed-rank test was conducted to determine whether, at each accuracy level, the median values of the differences between the MSEs or MAEs of the fusion forecasts and those of the naïve forecasts significantly deviated from zero. The MSEs and MAEs of the naïve and fusion forecasts and the results of the Wilcoxon tests are summarized in Table 1. Notably, the two p values derived when $ACC^{\text{movement}}=0.55$ are equal to zero, indicating a statistically significant enhancement of the fusion forecast over the naïve forecast, confirming that the observed improvements are not due to chance and underscoring the effectiveness of the fusion approach even at relatively low levels of $ACC^{\text{movement}}$. As $ACC^{\text{movement}}$ increases, the disparity between the naïve and fusion forecasts increases significantly.
@@
Prediction MSE MAE
Range p value Range p value
Naïve 1.044 - 0.821 -
Acc = 0.50, c = 0 1.044 - 0.821 -
Acc = 0.55, c = 0.1 1.037 $\pm$ 0.011 0.0000 0.816 $\pm$ 0.005 0.0000
Acc = 0.60, c = 0.2 1.017 $\pm$ 0.026 0.0000 0.802 $\pm$ 0.012 0.0000
Acc = 0.65, c = 0.3 0.982 $\pm$ 0.039 0.0000 0.775 $\pm$ 0.018 0.0000
Acc = 0.70, c = 0.4 0.939 $\pm$ 0.044 0.0000 0.745 $\pm$ 0.019 0.0000
Acc = 0.75, c = 0.5 0.885 $\pm$ 0.052 0.0000 0.712 $\pm$ 0.022 0.0000
Acc = 0.80, c = 0.6 0.794 $\pm$ 0.051 0.0000 0.664 $\pm$ 0.022 0.0000
Acc = 0.85, c = 0.7 0.716 $\pm$ 0.056 0.0000 0.623 $\pm$ 0.022 0.0000
Acc = 0.90, c = 0.8 0.618 $\pm$ 0.055 0.0000 0.578 $\pm$ 0.023 0.0000
Acc = 0.95, c = 0.9 0.493 $\pm$ 0.042 0.0000 0.526 $\pm$ 0.018 0.0000
Acc = 1, c = 1 0.371 - 0.482 -
Figure 3 illustrates the fusion results that are closest to the mean performance at each level of $ACC^{\text{movement}}$. When $ACC^{\text{movement}}$ slightly exceeds 0.50, such as in the second plot where $ACC^{\text{movement}}=0.55$, the fusion forecast shows only minor deviations from the naïve forecast and still exhibits a noticeable “shift” from the previous actual values. As $ACC^{\text{movement}}$ increases, this “shift” progressively diminishes and vanishes when $ACC^{\text{movement}}=1$.
Figure 4 illustrates the performance of the fusion forecasts across various accuracy levels using different fixed weights. As shown in the first four plots, under a smaller value of $c$, such as 0.01 or 0.1, the fusion forecast generally outperforms the naïve forecast, albeit at the expense of potential gains from higher $c$. Conversely, setting a larger $c$ could lead to inferior performance to the naïve forecast if the actual $ACC^{\text{movement}}$ is poor.
@@ 4.2 Experiment 2: Real-world Data
Table 3 presents the augmented Dickey–Fuller (ADF) test results for the two real-world time series. The ADF test can identify the presence of a unit root, assessing whether the S&P 500 and BTC price series exhibit random walk characteristics. For the S&P 500, the ADF statistic was 0.416927, with a p value of 0.982117. Given the critical values at the 1% (-3.433), 5% (-2.863), and 10% (-2.567) significance levels, the ADF statistic does not surpass these thresholds, and the high p value further supports the nonrejection of the null hypothesis. This suggests that the S&P 500 series likely follows a random walk, in which future values primarily reflect random fluctuations from previous values. Similarly, the BTC price series showed an ADF statistic of -1.449327 and a p value of 0.558413, with critical values not exceeded. This result indicates that the BTC price series is nonstationary and may also follow a random walk.
@@
Although the accuracy of financial time series movement forecasts can range between 0.55 and 0.8, this study opted to not employ complex models that integrate external data sources, which are resource intensive. Instead, this study utilized the nu-support vector machine (nu-SVM), a reliable binary classification model with the proportion of support vectors tuned through a parameter nu Bishop and Nasrabadi (2006). This choice allows us to focus on testing the robustness of the fusion method under real movement prediction with moderate accuracy rather than pursuing higher accuracy through more sophisticated models, as previous simulations have adequately demonstrated the performance of the fusion method across different accuracy levels. In addition, the details regarding the feature extraction process, which leverages a range of technical indicators and market statistics as model inputs, are outlined in Table LABEL:table:A1 in the Appendix. The nu-SVM prediction accuracy on the validation set was used as an estimation of $ACC^{\text{movement}}$ for setting the value of $c$ based on (11).
@@
Figure 6 and Figure 7 present the forecasts of different methods for the S&P 500 and BTC series, respectively. Because the movement prediction accuracy on the validation set is approximately 0.55 for both time series, the weight $c$ for the fusion method is set to 0.1. Consistent with the theoretical analysis presented in Section 3.2, this small weight $c$ leads to a small discrepancy between the naïve and fusion forecasts. Complex models such as MLP and LSTM differ significantly from fusion, naïve, and ARIMA forecasts because they presuppose a mapping from historical sequences to future values, which should depend solely on the current value with a random increment. The assumed patterns and dependencies in the data obtained by these models may not exist in true random walks.
Figure 6: The S&P 500 forecast results of different methods.
Figure 7: The BTC forecast results of different methods.
Table 5 summarizes the model performances in terms of the MSE and MAE values. The fusion method demonstrated the best overall performance among the models tested, achieving lower MSE and MAE values than did the naïve, ARIMA, and linear regression methods for both prediction tasks. Specifically, the fusion method yielded an MSE of 490.91 and an MAE of 16.19 for the S&P 500 prediction and an MSE of 449611.83 with an MAE of 431.94 for the BTC prediction, suggesting that the proposed method has a robust ability to make forecasts with reduced errors compared to other models. In contrast, both the MLP and LSTM models underperformed compared with the more traditional approaches, showing particularly high forecasting errors for both prediction tasks.
@@
MSE MAE MSE MAE
Fusion 490.912977 16.187902 449611.828348 431.940881
Naïve 499.378612 16.277100 456875.143422 433.766817
ARIMA 499.739638 16.338302 456466.453451 431.050006
Linear Regression 497.368764 16.236584 486329.417207 463.574310
MLP 775.427551 21.615049 486248.250000 496.776367
LSTM 690.336853 21.148510 503120.968750 457.213593
@@ 5 Discussion
This section discusses the implications of the findings based on theoretical analysis and experimental results and the potential limitations of this study. The empirical results show that when the $ACC^{\text{movement}}$ is approximately 0.55, the proposed fusion method outperforms the naïve forecast, which is consistent with the theoretical expectation that a movement prediction accuracy above 0.5 is necessary. Interestingly, the movement prediction accuracy of financial time series can reach as high as 0.8 in specific scenarios Bustos and Pomares-Quimbaya (2020), indicating that high movement prediction for real-world random walks is possible and can potentially lead to fusion results better than the examples shown in this study. Nonetheless, in real forecasting scenarios, $ACC^{\text{movement}}$ can only be estimated through model performance on the training or validation set, and such estimations can overestimate or underestimate the true $ACC^{\text{movement}}$. Whether $ACC^{\text{movement}}$ on the test set can consistently meet the accuracy threshold for improvement remains a challenge.
From the decision fusion perspective, integrating diverse predictions is advantageous because each model contributes unique insights and enhances predictive capabilities Bishop and Nasrabadi (2006). Previously, effectively combining predictions from various models that produce different types of data, such as binary and continuous outputs, has presented a significant challenge. The fusion method proposed in this study is an effective approach for converting binary classification into continuous values, thereby enabling the fusion of diverse data types. This conversion facilitates an efficient integration process, allowing for the comprehensive use of available forecasts to improve decision making in real forecasting scenarios.
Notably, the dataset used for experiments may not be large enough to meet the assumptions made in the theoretical analysis. In addition, while the definition of a random walk adopted in this study assumes i.i.d. increments for each time step, real-world data can be drifted random walks and are subject to measurement errors. All these factors could explain the discrepancies between the theoretical and empirical fusion results. Moreover, the random increments are supposed to be the same for all future time steps, which may not always provide a reliable estimation owing to the fluctuating nature of real-world data.
@@ 6 Conclusion
This study introduced a novel fusion method specifically designed for point forecasting in univariate random walks. Key findings reveal that the fusion method can outperform naïve forecasts with moderate movement prediction accuracies, such as 0.55. This improvement is achieved without consuming large amounts of computational resources and significantly exceeds that of many traditional forecasting methods, which often fail to surpass naïve forecasts. The implications of these results are profound, suggesting that the fusion method offers a potent approach for enhancing the forecasting of random walk series by effectively utilizing movement predictions.
@@
There are several avenues for further research that build upon the foundations laid by this study. First, although this research focuses on random walk forecasting, future work could expand the forecasting philosophy presented in this paper to encompass other types of time series data, potentially broadening its applicability. Second, given the dependency of the proposed method on movement predictions, which are still challenging in various areas, there is a pressing need to develop more advanced movement prediction techniques for random walk series. Such advancements, along with the fusion method proposed in this study, could contribute more to real random walk forecasting tasks.
Moreover, the current model utilizes a fixed increment for simplicity; however, adapting this increment based on the availability of new data may provide more dynamic and accurate forecasts. Investigating whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research. This adaptive approach could lead to more nuanced and responsive forecasting models, potentially transforming predictive analytics in various areas.